please help me
using the giving information write an equation of the line in slope intercept form find one perpendicular and one parallel for each problem
#1:(3,4) (1,3)

Answer by scott8148(6628)

the slope is the amount that the line rises (or falls) over some horizontal distance

[change in y] / [change in x]

m = (3 - 4) / (1 - 3) = -1 / -2 = 1/2

y = m x + b ___ 4 = 1/2 (3) + b ___ 5/2 = b

y = 1/2 x + 5/2

a perpendicular line will have a slope that is the negative reciprocal of the given slope

perpendicular ___ y = -2 x + 5/2

a parallel line will have the same slope but a different y-intercept (b value)

parallel ___ y = 1/2 x + 3